OpenStudy (aonz):

Help please! :) Show that the line x+y+2 = 0 is a tangent to y = x^3 −4x, and find the point of contact. [Hint: Find the equations of the tangents parallel to x+y +2 = 0, and show that one of them is this very line.]

hartnn (hartnn):

do you know how to find slope of tangent to y=x^3-4x ?

OpenStudy (aonz):

derive it 3x^2 - 4

hartnn (hartnn):

yeah, so the tangent parallel to x+y+2= 0 will have the SAME slope as that of x+y+2=0 so whats the slope of x+y+2=0 ??

OpenStudy (aonz):

slope of that is -1?

hartnn (hartnn):

can u find ?

OpenStudy (aonz):

so -1 = 3x^2 - 4?

OpenStudy (aonz):

ok thank you i think i can do it now

hartnn (hartnn):

oh great! welcome ^_^